# Debt, Growth, and the Carbon Lock-In

### v1

@@ B.5 Optimal leverage and carbon emissions
For strategy B, we get that the average growth rate is increasing with the leverage $\sim\left<\log\left(\gamma_{t}+L-1\right)\right$>, while the survival probability has a non monotonous complex behaviour as a function of leverage. We can build an optimization function between growth rate and the survival probability
$J(\beta,L,t_{p})=\beta\left<\log\left(\gamma_{t}+L-1\right)\right>+(1-\beta)P_{S}^{B}(t_{p})$ (73)
@@
Typically for the range of values corresponding to real data, we obtain the behaviour of Fig. 13. While the survival probability first decreases with leverage then plateaus or increases slowly, the growth rate keeps increasing with leverage. Therefore, the optimization function first decreases, reaches a minimum, and then increases again (the increase in growth rate compensates the plateau in survival probability). From this, we deduce that high leverages could be considered as optimal. However this optimization function does not account for carbon emissions resulting from a stronger growth rate. This suggests that the ”optimal” strategy with credit in a carbon intensive economy leads to stronger cumulative emissions than without credit.
@@
Figure 13: Optimization function for the leverage with $\beta=0.8,\rho=2,\left<\log(\gamma)\right>=0.7$ and for different values of $t_{p}$.
B.6 Optimal payback time based on solvency and short term growth rate
Let us start by reminding the key parameters in this version of the model for solvency:
  • $W$ is the long-term average growth rate, provided there was no bankruptcy; otherwise, this is 0.
  • $P_{S}(t_{p})$ is the probability of solvency after payback $t_{p}$.
  • $\sigma_{X}$ is the standard deviation of the long-term growth rate. It encapsulates the fluctuations of the system. We highlight that this is also the standard deviation of $m=W-\ln(\rho)$.
  • $WP_{S}(t_{p})$ is the effective long-term growth rate, conditioned on solvency after pay-back (otherwise the growth rate is 0).
  • The average growth rate of the capital relative to the initial investment $LC_{\text{ini}}$ just after the debt has been paid back, given there was no bankruptcy, $w(t_{p})$. It represents a short term growth rate.
The gambler wants to maximize the average growth rate $W$, the solvency probability $P_{S}(t_{p})$, minimize fluctuations $\sigma_{X}$ and maximize the average relative earnings after payback, keeping in mind that those parameters are interdependent.
We can first form the same optimization function as in [23]:
$J_{1}(\alpha)=\alpha W-(1-\alpha)\sigma_{X},$ (74)
where the value $\alpha$ sets the strategy. For $\alpha=1$ fluctuations (proxy for the risk) are ignored and only the growth rate is maximized, which corresponds to Kelly’s optimal strategy, while for $\alpha=0$ only fluctuations are minimized. Safe strategies correspond the strategies where fluctuations are reduced, even at the cost of reducing the growth rate.
If we optimize this with respect to the betting strategy $b_{x}$, under the constraint that $\sum_{i}b_{i}=1$ (Lagrange multiplier $\mu$), we get:
$\frac{\partial J_{1}}{\partial b_{i}}=$ $\alpha\frac{p_{i}}{b_{i}}-(1-\alpha)\frac{p_{i}}{\sigma_{X}b_{i}}\left(\ln(\frac{b_{i}}{r_{i}})-W\right)-\mu$ (75)
We deduce that:
$\mu b_{i}=\alpha p_{i}-\frac{1-\alpha}{\sigma_{X}}p_{i}\left(\ln\left(\frac{b_{i}}{r_{i}}\right)-W\right)$ (76)
Therefore, imposing that $\sum_{i}b_{i}=1$, we get a set of equations:
$\mu$ $=\alpha$ (77)
$b_{i}$ $=p_{i}\left(1-\frac{1-\alpha}{\alpha\sigma_{X}}\left(\ln\left(\frac{b_{i}}{r_{i}}\right)-W\right)\right)$ (78)
Thus we recover a correction from Kelly’s optimum $b_{i}=p_{i}$, depending on $\alpha,\sigma_{X}$ and $\boldsymbol{r}$. In this case, the bet should be modified particularly for horses corresponding to a value of $\ln(b_{i}/r_{i})$ significantly different from their average $W$. In particular, for a distribution of $b_{i}/r_{i}$ with fat tails.
The payback time is a balance between the need to accumulate as much as possible before payback and the risk of waiting too long before paying back, particularly with high interest rates. Indeed, if the interest rate is significantly smaller than the average earnings, the best strategy is to wait as long as possible to repay. Yet, if the interest rate is close to the average earnings, waiting too long may become risky due to fluctuations or simply because debts accumulate.
Strategy A
We observe that the Gaussian approximation seems to describe accurately the discrete case. Therefore we propose to use this expression of $P_{S}(t_{p})$ to optimize payback time with respect to solvency probability.
On another hand we can optimize the short term growth rate $w(t_{p})$ with respect to payback time $t_{p}$.
in a second optimization function with a new objective function $J_{2}(\beta)$, that interpolates between maximizing the solvency probability and maximizing short term growth rate relative to $\ln(E)$. We find
$J_{2}(\beta)=\beta\frac{1}{t_{p}\ln(E)}\ln\left(E^{t_{p}}-(1-\frac{1}{L})\rho^{t_{p}}\right)+\frac{1}{2}(1-\beta)\left(1+\text{erf}(\frac{t_{p}m-\ln\left(1-\frac{1}{L}\right)}{\sqrt{2t_{p}}\sigma_{X}})\right).$ (79)
The landscape of this optimization function is shown in Fig.14. Then, using $m=W-\ln(\rho)$, we are able to express:
$\begin{split}\frac{\partial J_{2}}{\partial t_{p}}=&\frac{\beta}{t_{p}^{2}\ln(E)}\left[t_{p}\frac{\ln(E)E^{t_{p}}-(1-\frac{1}{L})\ln(\rho)\rho^{t_{p}}}{E^{t_{p}}-(1-\frac{1}{L})\rho^{t_{p}}}-\ln\left(E^{t_{p}}-(1-\frac{1}{L})\rho^{t_{p}}\right)\right]\\ &+(1-\beta)\frac{W-\ln(\rho)+\frac{\ln(1-\frac{1}{L})}{t_{p}}}{2\sigma_{X}\sqrt{2\pi t_{p}}}\exp\left(-\frac{\left(t_{p}(W-\ln(\rho))-\ln\left(1-\frac{1}{L}\right)\right)^{2}}{2t_{p}\sigma_{X}^{2}}\right)\end{split}$ (80)
Figure 14: Landscape of the optimization function for two different sets of horses with $\beta=0.5$ in the limit of large leverage $L=15$.
Let us start with the limit of large $L$, we have
$\begin{split}\frac{\partial J_{2}}{\partial t_{p}}=&\frac{\beta}{t_{p}^{2}\ln(E)}\left[t_{p}\frac{\ln(E)E^{t_{p}}-\ln(\rho)\rho^{t_{p}}}{E^{t_{p}}-\rho^{t_{p}}}-\ln\left(E^{t_{p}}-\rho^{t_{p}}\right)\right]+(1-\beta)\frac{W-\ln(\rho)}{2\sigma_{X}\sqrt{2\pi t_{p}}}\exp\left(t_{p}\frac{\left(W-\ln(\rho)\right)^{2}}{2\sigma_{X}^{2}}\right)\end{split}$ (81)
We remind that the concavity of $\ln$ ensures that $W\leq\ln(E)$. We can compare three regions of the space $\left(\ln(\rho),\ln(E)\right)$ : $\ln(E)>W>\ln(\rho)$, $\ln(E)>\ln(\rho)>W$ and $\ln(\rho)>\ln(E)>W$ in the limit of large $L$
  • If $\ln(E)>W>\ln(\rho)$:
    $\frac{t_{p}^{3/2}(E^{t_{p}}-\rho^{t_{p}})\exp\left(\left(\left(\frac{(W-\ln(\rho))^{2}}{2\sigma_{X}^{2}}\right)-\ln(\frac{\rho}{E})\right)t_{p}\right)}{(\ln(E^{t_{p}})E^{t_{p}}-\ln(\rho^{t_{p}})\rho^{t_{p}})-(E^{t_{p}}-\rho^{t_{p}})\ln\left(E^{t_{p}}-\rho^{t_{p}}\right)}$ $=-\frac{\beta}{(1-\beta)\ln(E)}\frac{2\sqrt{2\pi}\sigma_{X}}{W-\ln(\rho)}<0,$ (82)
    this has no finite solution. Indeed we can show that $x\mapsto(x\ln(x)-\rho^{t_{p}}\ln(\rho^{t_{p}}))-(x-\rho^{t_{p}})\ln(x-\rho^{t_{p}})$ is an increasing function and thus $(\ln(E^{t_{p}})E^{t_{p}}-\ln(\rho^{t_{p}})\rho^{t_{p}})-(E^{t_{p}}-\rho^{t_{p}})\ln\left(E^{t_{p}}-\rho^{t_{p}}\right)\geq 0$. Considering that $J_{2}(\alpha,p_{x},b_{x},r_{x},t_{p})$ is increasing with $t_{p}$, it will be maximum for $t_{p}\to\infty$.
  • If $\ln(E)>\ln(\rho)>W$, then:
    $\frac{t_{p}^{3/2}(E^{t_{p}}-\rho^{t_{p}})\exp\left(\left(\left(\frac{(W-\ln(\rho))^{2}}{2\sigma_{X}^{2}}\right)-\ln(\frac{\rho}{E})\right)t_{p}\right)}{(\ln(E^{t_{p}})E^{t_{p}}-\ln(\rho^{t_{p}})\rho^{t_{p}})-(E^{t_{p}}-\rho^{t_{p}})\ln\left(E^{t_{p}}-\rho^{t_{p}}\right)}$ $=-\frac{\beta}{(1-\beta)\ln(E)}\frac{2\sqrt{2\pi}\sigma_{X}}{W-\ln(\rho)}>0,$ (83)
    this has finite solutions a priori.
  • If $\ln(\rho)>\ln(E)>W$, $J_{2}$, $w(t_{p})$ is not defined, we thus consider that the optimality is only set by maximizing the solvency probability. The optimum is thus $t_{p}\to 0$.
Now, when $L$ is small, new inequalities become decisive
$\begin{split}&\ln(\rho)<\ln(E)+\frac{1}{t_{p}}\ln(1+\frac{1}{L})\\ &\ln(\rho)<W+\frac{1}{t_{p}}\ln(1-\frac{1}{L}),\end{split}$ (84)
The first is the condition to have a well-defined short-term growth rate, and the second sets the transition to a solvency probability increasing with $t_{p}$ and a solvency probability decreasing with $t_{p}$. The first inequality means that the short-term growth rate can be defined even for $\ln(\rho)>\ln(E)$ (at small $t_{p}$). The second inequality means that the solvency probability can be a decreasing function of $t_{p}$ even for $W>\ln(\rho)$.
Figure 15: Optimal payback time depending on the conditions and the importance given to solvency probability. We observe that when we increase $\beta$, we grant less importance to solvency and waiting for longer becomes favorable. In addition, when increasing leverage $L$, we increase the short term growth rate, and we decrease solvency probability. Yet solvency probability becomes considerably smaller for large times, while the difference is smaller at early times. Therefore, increasing the leverage at fixed $\beta$ tends to increase optimal payback time. The region $E<\rho$ is never favorable.
We see on Fig.15 that the optimal payback time depends on $\rho$, $E$, $L$ and $\beta$. $\beta$ quantifies the relative importance given to short growth rate versus solvency probability. As $\beta$ increases, the gambler becomes less careful and wants to optimize the short-term growth rate rather than the solvency probability; in this case, the optimal payback time increases. As $L$ increases, the short-term growth rate also increases, and the solvency probability typically decreases, yet more significantly at long times. This tends to increase the optimal payback time. There are two main transitions in all diagrams, located around $E=\rho$ and $W=\ln(\rho)$ (shifted for small leverages), corresponding respectively to the transition where short-term growth rate starts being undefined at long times and to the transition where solvency probability starts decreasing with time.
B.7 Growth rate-solvency probability trade-off
The long-term growth rate is $W$ if the gambler survives to payback, and 0 otherwise. We can thus introduce an effective growth rate $WP_{S}(t_{p})$. We obtain the different dependencies of the parameters with respect to one another on Fig.16. We recover the results of [23] on Fig.16, where there is a trade-off branch where $\sigma_{X}$ is increasing with $WP_{S}(t_{p})$ (the effective growth rate). Kelly’s strategy corresponds to the maximal effective growth rate. We see on Fig.16 that $P_{S}(t_{p})$ is decreasing with $\sigma_{X}$ (increasing fluctuations means decreasing the probability of solvency), and that $w(t_{p})$ is increasing with $WP_{S}(t_{p})$ along the safe trajectory. Finally, we observe another trade-off in Fig.16, where $P_{S}(t_{p})$ decreases with $w(t_{p})$, meaning that ensuring solvency is done at the cost of reducing the short-term growth rate.
Figure 16: Different trade-offs between key parameters. In particular, increasing the effective growth rate goes along with increasing fluctuations. We also observe that increasing solvency probability goes along with decreasing short-term growth rate. Kelly’s strategy is shown as a square for each set of parameters, and safe strategies (varying $\alpha$) are shown as straight lines.
We highlight that for the safe strategy (full lines), fluctuations are smaller, the solvency probability is usually larger(except for large interest rates), the growth rate is smaller, and the post-payback earnings are smaller. The safe strategy corresponds to the optimization done in [23], with the parameter $\alpha$ varying.
@@ Appendix C Data
All data used in this work are publicly available for the different countries under study. In particular we get data for the US in [43, 45, 42, 44]. For France, we use [77, 78, 79, 80]. For Denmark, we use [81, 82, 83, 84]. For China, we use [85, 86, 87, 88].

### v2

@@ B.5 Optimal leverage and carbon emissions
For strategy B, we get that the average growth rate is increasing with the leverage $\sim\left<\log\left(\gamma_{t}+L-1\right)\right$>, while the survival probability has a non monotonous complex behavior as a function of leverage. Let us introduce an optimization function that interpolates between the growth rate and the survival probability:
$J(\beta,L,t_{p})=\beta\left<\log\left(\gamma_{t}+L-1\right)\right>+(1-\beta)P_{S}^{B}(t_{p}),$ (73)
where $\beta$ accounts for the relative weight of both contributions.
@@
Typically for the range of values corresponding to real data, we obtain the behavior of Fig. 12. While the survival probability first decreases with leverage then plateaus or increases slowly, the growth rate keeps increasing with leverage. Therefore, the optimization function first decreases, reaches a minimum, and then increases again (the increase in growth rate compensates the plateau in survival probability). From this, we deduce that high leverages could be considered as optimal. However this optimization function does not account for carbon emissions resulting from a stronger growth rate. This suggests that the ”optimal” strategy with credit in a carbon intensive economy leads to stronger cumulative emissions than without credit.
@@
Figure 12: Optimization function for the leverage with $\beta=0.8,\rho=2,\left<\log(\gamma)\right>=0.7$ and for different values of $t_{p}$. Decreasing $\beta$ leads to a stronger risk aversion, in which case the minimum of the optimization function is shifted to larger leverages. However the existence of a minimum does not depend on the value of $\beta$.
@@ Appendix C Data
All data used in this work are publicly available for the different countries under study. In particular, we get data for the US in [47, 49, 46, 48]. For France, we use [82, 83, 84, 85]. For Denmark, we use [86, 87, 88, 89]. For China, we use [90, 91, 92, 93].